OpenStudy (anonymous):

A circle is defined by the following equation: 3x^(2)+3y^(2)-24x+42y+183=0 What is its radius?

OpenStudy (anonymous):

i get \[\sqrt{126}\]

OpenStudy (anonymous):

let me try it slowly. \[3x^3+3y^3-24x+42y=-183\] \[x^2-8x + y^2 +14y=-61\] \[(x-4)^2+(y-7)^2=-61+16+49=4\] radius is 2

OpenStudy (anonymous):

should be \[(x-4)^2+(y+7)^2=4\] doesn't matter for radius, but center is (4,-7) not (4,7) hope the steps are clear: make the coefficient of the square terms 1 and then complete the square.

OpenStudy (anonymous):

where did you get the cubes from?